A Bibliography of Publications of Jörg Peters

Jörg Peters

Department of Computer & Information Science & Engineering

University of Florida Gainesville, FL 32611-6120

USA

Tel: +1 352 505-1576 FAX: +1 352 392-1220

E-mail: jorg@cise.ufl.edu

21 February 2025 Version 1.51

Abstract

This bibliography records publications of Jörg Peters.

Title word cross-reference

(3, 5) [Pet02a]. 2 [NKP16]. 3 [NKP16, Pet90a, PS04, PSZF06]. 4 [PS04]. A [KP15d]. C^{1} [KP15c, NKP16, NP16b, Pet90b, NP92, PS92b, Pet95b, PS92a, Pet95c]. C^{2} [KP05, KP09b, KP18a, KP18b, KP19a, MP09, Pet89, Pet02a, Pet96a, Pet95a]. C^{k} [GP15, Pet92b]. G [SP16]. G^{1} [ASC18, KP17, Pet18, Pet92a]. G^{2} [HPS12, KP11a, KP11b, KP15a, KP16]. G^{k} [GP15]. m [PS92a]. A_{n}^{∗} [KP10]. T [KP19b].

-constructions [GP15]. -continuous [GP15]. -direction [PS04]. -gons [KP19b]. -splines [SP16]. -surface [Pet95c]. -variate [PS92a].

1032 [PLK23]. 1996 [FBBD98].

3rd [FBBD98].

4-direction [NP92]. 4th [BBP^{+}08].

7-direction [Pet96a, Pet95a]. 783 [Pet98c].

8th [GHPW12].

’91 [MK91].

accurate [YBP14]. Advances [BBP^{+}08]. affine [PR98a]. aided [DLS94, Sap94]. Alberta [MK91]. Algorithm [OPB^{+}18, Pet98c, PLK23]. Algorithms [PR98b]. almost [KP19a]. always [GP15]. analogues [NP92]. Analysis [PR98b, MP23, RP06]. animated [YBP14]. application [OPR06]. Applications [Mul96]. Approximate [Pet94c]. approximating [KP12a]. approximation [BS92]. arbitrary [ASC18, Pet94b, Pet97a, Pet18].

Assembling [KP09a]. Atlasing [OPB^{+}18]. Attraction [KP23]. augmented [KP15c].

B [Pet97a, PR98b, Pet98c, Pet98a]. B-Spline [PR98b]. B-splines [Pet97a, Pet98c, Pet98a]. based [KP14, Mul96, PW97b, PW03]. basic [KP11b, KP12b, KP13a, PW97a, PR04]. BB [KP09c]. BB-form [KP09c]. be [KP15d]. Bernstein [Pet94c]. better [Pet03]. between [NPLxx, SPZ10]. Bézier [ASC18, NPLxx, Pet94c, Pet18, GP95]. Bi [MP09, PLK23, ASC18, KP11b, KP15d, KP16, MKP08, Pet18]. Bi-3 [MP09]. bi-6 [KP16]. Bi-cubic [PLK23, ASC18, KP11b, KP15d, MKP08, Pet18]. Bicubic [KP07, Pet98b, KP16, Pet90b, Pet97a, Pet97b, Pet98c, Pet98a]. bicubics [Pet94b]. biquadratic [KP15c, KP15e]. biquadratics [Pet94b]. Biquartic [Pet95b]. Biquintic [KP15a]. Blending [PW97a, KP15e]. blends [KP15b, PW97b]. bodies [SPZ10]. Boehm [Pet95d]. Book [Pet95d]. Boston [War92]. Boundary [NP17, Pet90b]. bounded [PU01]. bounds [NPLxx, PU01]. Box [KEP08, PW97b, KP09c, KP10, KP11c, KP24, Pet96a, Pet95a, PW97a]. Box-spline [PW97b]. box-splines [KP09c, KP10, KP11c]. Brunel [Mul96]. built [Pet96a, Pet95a].

C [Pet97a]. CAGD [HP94, PH94]. Calgary [MK91]. Can [KP15d]. case [KPR04]. Centered [KEP08]. Changing [Pet12]. characterization [KPR04, Pet94a, PR04]. Class [KP18a, KP15d]. classes [PR98a]. classification [Pet90c]. code [Pet97a]. Collision [WP04]. Combinatorial [SZP10]. Combining [KP18a, PS04, KP12b]. compatible [HPS12]. compendium [KP24]. completion [KP16]. complexity [PF10]. composite [Pet92a]. computer [DLS94, Sap94, War92]. computer-aided [Sap94]. Computing [BBP^{+}08, GOMP98, PN97, PU01]. Concepts [BPP95, Pet95d]. Conference [GM97, Mul96]. Configuration [OPB^{+}18]. Conflicting [SZP10]. connecting [Pet94a]. connectivity [MKP08]. Constrained [OPB^{+}18]. Constraint [SZP10, PSZF06]. Constraints [HP94, PH94, YP18]. constructed [Pet94b]. Construction [MNP08]. constructions [GP15, MKP08]. continuity [Pet02b]. continuous [GP15, KP09a, Pet96b]. contracting [KP18b]. Control [PLK23, MP23, NPLxx, Pet96d]. control-net [MP23]. convex [Pet96c]. convexity [GP95]. CSG [PW97b]. Cubic [KEP08, ASC18, KP11b, KP15d, MKP08, Pet90b, Pet90e, Pet96c, Pet18, PLK23]. cubics [FP96, PS92b]. Curvature [KP12a, KP13a, Pet96b, KP09a, PU00, PU01]. Curvature-sensitive [KP13a]. Curve [HPS12, FP96, KP12a, Pet96c]. curved [Pet04, YBP14]. Curves [DLS98, Pet89, War92, KP13a, Pet91b, Sap94].

D [Pet90a, PSZF06, NKP16]. Dagstuhl [FBBD98, GHPW12, HFPW09]. data [Pet90a]. December [BBP^{+}08]. Degree [NP17, KP14, Pet02a]. dependent [OPR06]. derived [Pet14]. Design [BPP95, DLS94, FP96, KP11b, KP13a, Sap94, Pet95d]. Designing [Sap94]. Detection [WP04]. Different [GTP98]. dimensional [GP95, NP92]. direction [NP92, Pet96a, Pet95a, PS04]. Discontinuous [NP16a, NP17, Pet15]. Discrete [PRR05]. Distance [OPB^{+}18, WP04, NPLxx]. dominant [KP19b]. double [FP96]. double-Tschirnhaus [FP96]. Dundee [GM97].

Easy [Pet95e]. Editorial [HFPW09].

effectively [KP18b]. Efficient [OPB^{+}18, YBP14]. elements [GP15, NKP16, NP16b]. Elimination [PSZF06]. enclose [Pet91a]. Enclosed [PN97, GOMP98]. enclosures [LP01a, PW03]. envelopes [LP01b]. Equations [PRR05, OPR06]. equivalence [PR98a]. Evaluation [Pet94c, KP09c]. everywhere [KP19a]. Exchange [GTP98]. Explicit [NP17].

Face [KEP08]. Face-Centered [KEP08]. faces [KP22a]. Facets [MNP08]. Fair [KP19a, Sap94]. fairness [Pet03]. Fast [KP09c, MNP08, PRR05]. Fields [WP04]. Filters [NP16a, NP17, Pet15]. finite [KP18b, NKP16]. Fitting [Pet90a]. Floorplanning [HAPR94]. Foreword [BMP22]. Form [Pet94c, KP12b, KP17, KP19a, KP09c, Pet92b, Pet93, Pet97a, Pet98c, Pet98a, Pet02a]. Free [KP12b, KP17, KP19a, Pet93, Pet97a, Pet98c, Pet98a, Pet02a]. Free-form [KP12b, KP17, KP19a, Pet93, Pet97a, Pet98c, Pet98a, Pet02a]. functionals [KP15a]. Functions [PW06, KP17].

Galerkin [NP16a, NP17, Pet15]. Gaussian [PU00]. General [Pet15]. Generalized [Pet89, PW06, OPR06]. Generalizing [PR98b, Pet93]. generically [PSZF06]. geodesic [YP18]. Geometric [BPP95, FBBD98, GHPW12, HP94, PH94, Pet95d, Pet02b, SZP10, DLS94, HFPW09, PSZF06, Sap94, Pet95d]. geometry [Pet04]. Germany [FBBD98]. gons [KP19b]. Good [KP18a]. GPU [SJP05]. Graphics [MK91, War92]. Guided [KP18a, KP05].

H [Pet95d]. held [Mul96]. Hermite [Pet89]. higher [GP95, NP92]. higher-dimensional [NP92]. house [GP98].

identity [Pet94a]. II [DLS98]. III [DLS94, War92]. implicit [PW97a]. Improved [KP15b, KP22a]. Improving [Pet92a]. incidences [SPZ10]. Independence [PW06]. Institute [Mul96]. Integration [HAPR94]. Interactive [Gon97, HAPR94]. Interface [MK91, OPR06]. Interference [WP04]. International [BBP^{+}08, BMP22]. interpolants [KP14]. Interpolation [Pet89, Pet96c, Pet90b, Pet90c, Pet90e, Pet91b, NP92, PS92b, PS92a, YP18]. interpolatory [KP13b, KP14]. Intervals [OPB^{+}18]. irregular [NP16b, Pet93, Pet95b, Pet96b, Pet96d]. isogeometric [GP15]. Issue [GHPW12, BMP22, HFPW09]. ISVC [BBP^{+}08]. Iterative [PRR05].

Joining [Pet92b]. joins [Pet92a]. junctions [KPP17]. June [MK91].

kernel [SJP05]. knot [Pet13].

Lattice [KEP08, KP10]. lattices [KP11c]. layout [NP16b, Pet97a]. Learning [HAPR94]. Least [NP17]. Least-Degree [NP17]. Lens [KP09b]. Lens-shaped [KP09b]. Lillehammer [DLS98]. Linear [PW06, LP01b, PW03, Pet04, PS15]. linearly [Pet90b, Pet97a, Pet98c, Pet98a]. linearly-trimmed [Pet97a, Pet98a]. linking [Pet04]. Local [Pet89, Pet90b, Pet90c, PW06, KP14]. Localized [KP22b]. locally [KP19b].

Made [Pet95e]. manifolds [NKP16]. maps [Pet94a]. Massachusetts [War92]. Matched [GP15]. Mathematical [DLS94, DLS98]. Mathematics [GM97, Mul96]. max [PW03]. max-norm [PW03]. mean [PU00]. mesh [Pet90e, Pet91b, NP92]. Meshes [MNP08, KP19b, NNP07, Pet93, Pet95b, Pet96b, Pet96d, PF10]. Messages [GTP98]. method [Pet90d]. Methods [DLS98, BS92, DLS94]. Mid [Pet04].

Mid-structures [Pet04]. Minimal [KP16, KP14]. Modeling [BMP22, Gon97, HFPW09, Pet97a, Sap94, GHPW12]. modelling [FBBD98]. moments [GOMP98]. multi [KP15b, KP15e, KP22a, Pet03]. multi-sided [KP15e, KP22a, Pet03]. multi-surface [KP15b]. multiprocessor [Pet90d]. multisided [KP05]. Multistrategy [HAPR94]. Multivariate [Pet94c, LP01a].

n [PR98a]. n-space [PR98a]. need [Pet03]. Nested [KP18a]. net [MP23]. Nets [PLK23, GP95]. network [Pet90d, YP18]. networks [HPS12]. Non [KP13b, KP14, NKP16]. non-tensor-product [NKP16]. Non-uniform [KP13b, KP14]. Nonuniform [NP16a]. norm [PW03]. normal [Pet90b]. November [War92]. Numerical [BS92]. NURBS [KP12b, Pet97a, Pet97b]. NV [BBP^{+}08].

objects [GOMP98]. optimality [PW03]. Optimized [LP01a, SPZ10]. organized [Mul96].

Pairs [MKP08]. papers [GHPW12]. Parallel [MNP08]. parameter [OPR06]. parametric [Pet90a, Pet91a]. parametrically [KP19a]. parametrization [SPZ10]. Parametrizing [Pet91a]. Part [BBP^{+}08]. Partitioned [Pet94c]. patch [Pet92a, Pet97a]. patch-layout [Pet97a]. Patches [Pet98b, ASC18, KP05, KP09a, Pet90e, Pet92b, Pet03, Pet18]. Patching [Pet01]. Pcp2Nurb [Pet98c, Pet98a]. Pent [MNP08]. piece [NPLxx]. pieces [LP01a]. piecewise [GOMP98, PW03, Pet04]. pixel [YBP14]. pixel-accurate [YBP14]. Platonic [PK98]. Point [KP15c, OPB^{+}18, OPR06]. Point-augmented [KP15c]. Point-Sets [OPB^{+}18]. polar [KP07, MKP08, MP09]. polycube [SP16]. polygon [NPLxx]. Polygons [KP15d]. Polyhedra [Pet95e, Pet98b, Pet97b, PR97]. Polyhedral [MP23, PLK23, KP22b]. Polynomial [KP05, PS15, GOMP98, LP01a, NPLxx]. practical [KP24]. Prautzsch [Pet95d]. preconditioners [OPR06]. Preprocessors [SZP10]. principles [PR04]. problem [OPR06]. Proceedings [BBP^{+}08, GM97, MK91, Mul96]. product [NKP16].

Quad [MNP08, KP19b, NP16b, PF10]. quad-dominant [KP19b]. Quadratic [KP23, Pet93, PR98a]. Quadratic-Attraction [KP23]. quadrilateral [NNP07]. Quads [KP15d]. quality [Sap94]. quantitative [NPLxx]. Quartic [Pet89].

Rapidly [KP18b]. Rational

[KP11a, KP11b, KP05, PS15]. realtime [SJP05]. Reconciling [SZP10]. Reconstruction [KEP08]. Recursive [PN97]. Refinability [Pet14]. Refinable [KP17, KP19b, NP16b, SP16, LP01a]. Refined [Pet01]. Refinement [KP18a, KP22a]. regions [Pet96c]. regular [Pet14]. Regularly [Pet94c]. remeshing [KP22b]. rendering [YBP14]. Response [ASC18]. Review [Pet95d]. rigid [PSZF06, SPZ10]. root [KP11c]. roots [Pet94a]. rule [KP22a].

saddle [OPR06]. Scale [NP16a]. scheme [PR97]. schemes [KP12a]. Search [OPB^{+}18]. Section [KP11b]. segments [Pet96c]. selected [GHPW12]. seminar [GHPW12]. sensitive [KP13a]. September [GM97, Mul96]. sequences [Pet13]. Sets [OPB^{+}18, Pet96a, Pet95a]. Shape [BMP22, KPR04, KP18a, PR04, KP15b, Sap94]. shaped [KP09b]. shapes [KP11b, KP12b, PW97a]. Sharp [NPLxx].

Shift [NP16a]. sided [KP15e, KP22a, Pet03]. simplest [PR97]. Simplex [Pet94c, Pet90d]. simplices [GP95]. singularly [Pet91a]. Sizes [GTP98]. slefes [PW03]. SMI2022 [BMP22]. Smooth [FP96, KP15e, MNP08, Pet90e, Pet91b, Pet93, Pet01, KP19b, Pet90a, Pet90c, Pet91a, Pet92b, Pet97a, Pet98c, Pet98a, PF10]. Smoothing [Pet95e, Pet97b, Pet98b, PR97]. Smoothness [Pet03]. Solids [PN97]. solutions [Pet15]. Solvers [PRR05]. Space [Pet89, PR98a]. Spaces [OPB^{+}18]. Special [GHPW12, BMP22, HFPW09]. Spheroids [PK98]. Spline [KPP17, KEP08, Pet96d, PR98b, KP05, KP16, KP24, NP16b, Pet95b, Pet96a, Pet95a, Pet96b, PW97b, PF10, Pet15, PS15]. Splines [Gon97, Pet13, PLK23, KP11a, KP11b, KP12b, KP13a, KP13b, KP15e, KP22b, KP09c, KP10, KP11c, LP01b, MP23, Pet93, Pet95c, PW97a, Pet97a, Pet98c, Pet98a, Pet14, SP16]. Stability [PS92b, PS92a]. stable [KP09c]. stitched [ASC18, Pet18]. Stokes [OPR06, PRR05]. Structural [RP06]. structures [Pet04]. studies [KPR04]. Subdivision [KP18a, KP23, PN97, PR98b, PW06, PR08, WP04, GP95, KPR04, KP07, KP09b, KP12a, KP13b, KP14, KP15c, KP18b, MP09, NNP07, PR97, PU00, PU01, PS04, PR04, RP06, SJP05]. summary [RP06]. supporting [MKP08]. Surface [Gon97, KP15b, MKP08, Pet90b, Pet90c, Pet91a, Pet92a, Pet92b, Pet95c]. Surfaces [DLS98, KPP17, KP18a, MNP08, Pet94b, PN97, WP04, GOMP98, GM97, KPR04, KP05, KP09a, KP09b, KP15a, KP15c, KP15d, KP16, KP17, KP18b, KP19a, KP19b, Mul96, Pet90a, Pet93, Pet95b, Pet96a, Pet95a, Pet96b, Pet96d, Pet97a, PR98a, PU00, Pet02a, PR04, PR08, PF10, PS15, RP06, Sap94, War92, YBP14]. surfacing [HPS12, Pet98c, Pet98a].

Symmetric [KP10, KP11c]. Symposium [BBP^{+}08]. System [HAPR94]. Systems [SZP10, PSZF06, SPZ10].

T [KPP17]. T-junctions [KPP17]. tensor [NKP16]. Ternary [NNP07]. tessellations [Pet14]. theory [BS92]. Tight [LP01b]. Tool [HAPR94]. topological [GP98]. topology [ASC18, Pet94b, Pet18]. transitions [PS15]. Tri [MNP08]. Tri/Quad/Pent [MNP08]. triangular [KP09a]. triangulation [PS92b]. Triangulations [Pet01]. Trimmed [Pet98b, Pet97a, Pet97b, Pet98c, Pet98a]. trivariate [PW97a]. Tschirnhaus [FP96].

UK [GM97]. underlying [PS92b]. Uniform [OPR06, KP13b, KP14]. University [Mul96]. unsorted [Pet13]. USA [BBP^{+}08]. Using [Pet98b, Gon97, Pet92a, PW97a, Pet97b].

Variables [Pet12]. variate [PS92a].

varying [Pet90b]. Vegas [BBP^{+}08]. vertex [Pet92b, YP18]. vertices [Pet91a, PS92b]. VI [Mul96]. via [KP13b, KP15a, KP09c, NP16a]. VII [GM97]. vision [War92]. Visual [BBP^{+}08]. Volume [BS92]. Volumes [PN97].

W [Pet95d]. workshop [FBBD98].

XXX [Far97].

yield [GP15]. yields [KP18b].

zero [Pet96a, Pet95a].

References

Akleman:2018:RT

[ASC18] Ergun Akleman, Vinod Srinivasan, and Jianer Chen. Response to “On G^{1} stitched bi-cubic Bézier patches with arbitrary topology”. Computers and Graphics, 71(??):157–158, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic).  URL https://www.
sciencedirect.com/science/
article/pii/S0097849317301954.

See [Pet18].

Bebis:2008:AVCa

[BBP^{+}08] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Paolo Remagnino, Fatih Porikli, Jörg Peters, James Klosowski, Laura Arns, Yu Ka Chun, Theresa-Marie Rhyne, and Laura Monroe, editors. Advances in Visual Computing: 4th International Symposium, ISVC 2008, Las Vegas, NV, USA, December 1–3, 2008. Proceedings, Part I, volume 5358 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CODEN LNCSD9. ISBN 3-540-89638-4 (print), 3-540-89639-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/
content/978-3-540-89639-5.

Biasotti:2022:FSIc

[BMP22] Silvia Biasotti, Ramanathan Muthuganapathy, and Jörg Peters. Foreword to the special issue on shape modeling international 2022 (SMI2022). Computers and Graphics, 107(??): A6–A8, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0097849322001613.

Boehm:1995:GCG

[BPP95] W. Boehm, H. Prautzsch, and Jörg Peters. Geometric concepts for geometric design. SIAM Review, 37(3):473–??, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Braess:1992:NMA

[BS92] D. Braess and L. L. Schumaker, editors. Numerical methods in approximation theory, Volume 9, volume 105 of International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1992. ISBN 3-7643-2746-4 (Basel), 0-8176-2746-4 (Boston). xiv + 357 pp. LCCN QA221 .N84 1992.

Daehlen:1994:MMC

[DLS94] M. Dæhlen, T. Lyche, and L. L. Schumaker, editors. Mathematical methods in computer aided geometric design III. Vanderbilt University Press, Nashville, TN, USA, 1994. ISBN 0-8265-1268-2. LCCN QA565 .M37 1995.

Daehlen:1998:MMC

[DLS98] Morten Dæhlen, Tom Lyche, and Larry L. Schumaker, editors. Mathematical Methods for Curves and Surfaces II: Lillehammer, 1997. Vanderbilt University Press, Nashville, TN, USA, 1998. ISBN 0-8265-1315-8. LCCN QA565 .M372 1998.

Farin:1997:XX

[Far97] G. Farin, editor. XXX. Springer-Verlag, Berlin, 1997.

Farin:1998:GMD

[FBBD98] Gerald Farin, Hanspeter Bieri, Guido Brunnett, and Tony DeRose, editors. Geometric modelling. 3rd Dagstuhl workshop, Dagstuhl, Germany, 1996, volume 13 of Computing. Supplementum. Springer, Wien / New York, 1998. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998.

URL http://www.springer.at/
springer.py?Page=30&Key=340&
id=589&cat=3.

Farouki:1996:SCD

[FP96] R. Farouki and J. Peters.

Smooth curve design with double-Tschirnhaus cubics. Annals of Numerical Mathematics, 3(??), ???? 1996. ISSN 1021-2655.

Grandine:2012:SIS

[GHPW12] Thomas Grandine, Stefanie Hahmann, Jörg Peters, and Wenping Wang. Special issue of selected papers from the 8th Dagstuhl seminar on Geometric Modeling. Graphical models, 74 (6):291, November 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S1524070312000422.

Goodman:1997:MS

[GM97] Tim N. T. Goodman and Ralph R. Martin, editors. The Mathematics of surfaces VII: Proceedings of the Mathematics of Surfaces VII Conference, Dundee, UK, September 1996. Information Geometers, Winchester, UK, 1997. ISBN 1-874728-12-7.

Gonzalez-Ochoa:1998:CMO

[GOMP98] Carlos Gonzalez-Ochoa, Scott McCammon, and Jörg Peters. Computing moments of objects enclosed by piecewise polynomial surfaces. ACM Transactions on Graphics, 17(3):143–157, July 1998. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL http:/

                                        /www.acm.org/pubs/citations/
                                        journals/tog/1998-17-3/p143-
                                        gonzalez-ochoa/.

Gonzalez:1997:IMU

[Gon97] C. Gonzalez. Interactive modeling using surface splines. In 1er Encuentro de Computacion, ENC 97, page ?? Sociedad Mexicana de Ciencia de la Computacion, 1997. Queretaro, Mexico.

Goodman:1995:BNC

[GP95] T. N. T. Goodman and J. Peters. Bézier nets, convexity and subdivision on higher dimensional simplices. Computer-Aided Geometric Design, 12:53–65, 1995. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Gonzalez:1998:TH

[GP98] C. Gonzalez and J. Peters. The topological house, 1998. URL

http://www-sfb288.math.tu-
berlin.de/VideoMath/program/
page27.html. Videotape.

Groisser:2015:MCA

[GP15] David Groisser and Jörg Peters.

Matched G^{k}-constructions always yield C^{k}-continuous isogeometric elements. Computer-Aided Geometric Design, 34:67–72, 2015. CODEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Goldman:1998:EMD

[GTP98] A. Goldman, D. Trystram, and J. Peters. Exchange of messages of different sizes. Lecture Notes in Computer Science, 1457: 194–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrmann:1994:MLS

[HAPR94] J. Herrmann, R. Ackermann, J. Peters, and D. Reipa. A multistrategy learning system and its integration into an interactive floorplanning tool. Lecture Notes in Computer Science, 784: 138–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahmann:2009:ESI

[HFPW09] Stefanie Hahmann, Gerald Farin, Jörg Peters, and Wenping Wang. Editorial: Special issue on geometric modeling (Dagstuhl 2008). Computing, 86(2–3): 71–72, October 2009. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0010-485X&
volume=86&issue=2&spage=71.

Hoffmann:1994:GCC

[HP94] C. M. Hoffmann and J. Peters. Geometric constraints of CAGD. In Dæhlen et al. [DLS94], pages 369–376. ISBN 0-8265-1268-2. LCCN QA565 .M37 1995.

Hermann:2012:CNC

[HPS12] Thomas Hermann, Jörg Peters, and Tim Strotman. Curve networks compatible with G^{2} surfacing. Computer-Aided Geometric Design, 29(5):219–230, 2012. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Kim:2008:BSR

[KEP08] Minho Kim, Alireza Entezari, and Jörg Peters. Box spline reconstruction on the face-centered cubic lattice. IEEE Transactions on Visualization and Computer Graphics, 14(6):1523–1530, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Karciauskas:2005:PSS

[KP05] Kȩstutis Karčiauskas and Jörg Peters. Polynomial C^{2} spline surfaces guided by rational multisided patches. In Computational methods for algebraic spline surfaces, pages 119–134. Springer-Verlag Inc., New York, NY, USA, 2005.

Karciauskas:2007:BPS

[KP07] Kȩstutis Karčiauskas and Jörg Peters. Bicubic polar subdivision. ACM Transactions on Graphics, 26(4):14:1–14:6, October 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karciauskas:2009:ACC

[KP09a] Keştutis Karciauskas and Jörg Peters. Assembling curvature continuous surfaces from triangular patches. Computers and Graphics, 33(3):204–210, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0097849309000302.

Karciauskas:2009:LSS

[KP09b] Kestutis Karciauskas and Jörg Peters. Lens-shaped surfaces and C^{2} subdivision. Computing, 86(2– 3):171–183, October 2009. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0010-485X&
volume=86&issue=2&spage=171.

Kim:2009:FSE

[KP09c] Minho Kim and Jörg Peters. Fast and stable evaluation of boxsplines via the BB-form. Numerical Algorithms, 50(4):381– 399, April 2009. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

                                        http://www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=1017-1398&volume=50&issue=
                                        4&spage=381.

Kim:2010:SBS

[KP10] Minho Kim and Jörg Peters. Symmetric box-splines on the A_{n}^{∗} lattice. Journal of Approximation Theory, 162(9):1607–1630, 2010. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Karciauskas:2011:RS

[KP11a] Kȩstutis Karčiauskas and Jörg Peters. Rational G^{2} splines. Graphical models, 73(5):286– 295, September 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S152407031100021X.

Karciauskas:2011:SRB

[KP11b] Kȩstutis Karčiauskas and Jörg Peters. Section 1: Rational bicubic G^{2} splines for design with basic shapes. Computer Graphics Forum, 30(5):1389–1395, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2011:SBS

[KP11c] Minho Kim and Jörg Peters. Symmetric box-splines on root lattices. Journal of Computational and Applied Mathematics, 235(14):3972–3989, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Karciauskas:2012:CAC

[KP12a] Kȩstutis Karčiauskas and Jörg Peters. Curvature of approximating curve subdivision schemes. In Curves and surfaces, volume 6920 of Lecture Notes in Computer Science, pages 369–381. Springer-Verlag Inc., New York, NY, USA, 2012.

Karciauskas:2012:FFS

[KP12b] Keştutis Karciauskas and Jörg Peters. Free-form splines combining NURBS and basic shapes. Graphical models, 74(6):351– 360, November 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S1524070312000379.

Karciauskas:2013:CSS

[KP13a] Kȩstutis Karčiauskas and Jörg Peters. Curvature-sensitive splines and design with basic curves. Computer-Aided Design, 45(2): 415–423, 2013. CODEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (electronic).

Karciauskas:2013:NUI

[KP13b] Kȩstutis Karčiauskas and Jörg Peters. Non-uniform interpolatory subdivision via splines. Journal of Computational and Applied Mathematics, 240:31–41, 2013. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Karciauskas:2014:NUI

[KP14] Kȩstutis Karčiauskas and Jörg Peters. Non-uniform interpolatory subdivision based on local interpolants of minimal degree. In Mathematical methods for curves and surfaces, volume 8177 of Lecture Notes in Computer Science, pages 248–264. Springer-Verlag Inc., New York, NY, USA, 2014.

Karciauskas:2015:BSF

[KP15a] Kȩstutis Karčiauskas and Jörg Peters. Biquintic G^{2} surfaces via functionals. Computer-Aided Geometric Design, 33:17–29, 2015. CODEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Karciauskas:2015:ISM

[KP15b] Kȩstutis Karciauskas and Jörg Peters. Improved shape for multisurface blends. Graphical models, 82(??):87–98, November 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S1524070315000259.

Karciauskas:2015:PAB

[KP15c] Keştutis Karciauskas and Jörg Peters. Point-augmented biquadratic C^{1} subdivision surfaces. Graphical models, 77 (??):18–26, January 2015. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S152407031400054X.

Karciauskas:2015:QPC

[KP15d] KȨstutis Karčiauskas and Jörg Peters. Quads and polygons: Can bi-cubic surfaces be class A? Computer Graphics Forum, 34(5):229–238, August 2015. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karciauskas:2015:SMS

[KP15e] Kȩstutis Karciauskas and Jörg Peters. Smooth multi-sided blending of biquadratic splines. Computers and Graphics, 46(??): 172–185, February 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0097849314000922.

Karciauskas:2016:MBC

[KP16] Kȩstutis Karčiauskas and Jörg Peters. Minimal bi-6 G^{2} completion of bicubic spline surfaces. Computer-Aided Geometric Design, 41:10–22, 2016. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Karciauskas:2017:RFF

[KP17] Kȩstutis Karčiauskas and Jörg Peters. Refinable G^{1} functions on G^{1} free-form surfaces. Computer-Aided Geometric Design, 54:61– 73, 2017. CODEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Karciauskas:2018:NCG

[KP18a] Kȩstutis Karčiauskas and Jörg Peters. A new class of guided C^{2} subdivision surfaces combining good shape with nested refinement. Computer Graphics Forum, 37(6):84–95, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karciauskas:2018:RCS

[KP18b] Kȩstutis Karciauskas and Jörg Peters. Rapidly contracting subdivision yields finite, effectively C^{2} surfaces. Computers and Graphics, 74(??):182–190, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic).  URL https://www.
sciencedirect.com/science/
article/pii/S0097849318300839.

Karciauskas:2019:FFF

[KP19a] Kȩstutis Karciauskas and Jörg Peters. Fair free-form surfaces that are almost everywhere parametrically C^{2}. Journal of Computational and Applied Mathematics, 349(??):470–481, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042718304667.

Karciauskas:2019:RSS

[KP19b] Kȩstutis Karciauskas and Jörg Peters. Refinable smooth surfaces for locally quad-dominant meshes with T -gons. Computers and Graphics, 82(??):193–202, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0097849319300743.

Karciauskas:2022:IRR

[KP22a] Kȩstutis Karciauskas and Jörg Peters. An improved refinement rule for multi-sided faces. Computers and Graphics, 102(??): 370–379, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0097849321002193.

Karciauskas:2022:LRP

[KP22b] Kȩstutis Karciauskas and Jörg Peters. Localized remeshing for polyhedral splines. Computers and Graphics, 106(??): 58–65, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0097849322000905.

Karciauskas:2023:QAS

[KP23] K. Karčiauskas and J. Peters.

Quadratic-attraction subdivision. Computer Graphics Forum, 42(5): e14900:1–e14900:??, August 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2024:PBS

[KP24] Minho Kim and Jörg Peters. A practical box spline compendium. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300323005453.

Karciauskas:2017:JSS

[KPP17] Kȩstutis Karciauskas, Daniele Panozzo, and Jörg Peters. Tjunctions in spline surfaces. ACM Transactions on Graphics, 36(5): 170:1–170:??, October 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karciauskas:2004:SCS

[KPR04] K. Karčiauskas, Jörg Peters, and U. Reif. Shape characterization of subdivision surfaces—case studies. Computer-Aided Geometric Design, 21(6):601–614, 2004. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Lutterkort:2001:ORE

[LP01a] David Lutterkort and Jörg Peters. Optimized refinable enclosures of multivariate polynomial pieces. Computer-Aided Geometric Design, 18(9):851–863, 2001. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Lutterkort:2001:TLE

[LP01b] David Lutterkort and Jörg Peters. Tight linear envelopes for splines. Numerische Mathematik, 89(4):735–748, October 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

                                        //link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        1089004/10890735.htm;   http:
                                        //link.springer-ny.com/link/

service/journals/00211/papers/
1089004/10890735.pdf.

MacKay:1991:GIC

[MK91] S. MacKay and E. M. Kidd, editors. Graphics Interface ’91, Calgary, Alberta, 3–7 June 1991: proceedings. Canadian Information Processing Society, 243 College St, 5th Floor, Toronto, Ontario M5T 2Y1, Canada, 1991. ISBN 0-9695338-0-2. LCCN T385.P792 1991.

Myles:2008:PBC

[MKP08] Ashish Myles, K‘estutis Karčiauskas, and Jörg Peters. Pairs of bi-cubic surface constructions supporting polar connectivity. Computer-Aided Geometric Design, 25(8):621–630, 2008. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Myles:2008:SSF

[MNP08] A. Myles, T. Ni, and J. Peters.

Smooth surfaces: Fast parallel construction of smooth surfaces from meshes with tri/quad/pent facets. Computer Graphics Forum, 27(5):1365–1372, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Myles:2009:BCP

[MP09] Ashish Myles and Jörg Peters. Bi-3 C^{2} polar subdivision. ACM Transactions on Graphics, 28(3): 48:1–48:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mishra:2023:PCN

[MP23] Bhaskar Mishra and Jörg Peters. Polyhedral control-net splines for analysis. Computers and Mathematics with Applications, 151 (??):215–221, December 1, 2023. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0898122123004261.

Mullineux:1996:MS

[Mul96] Glen Mullineux, editor. The Mathematics of surfaces VI: based on the proceedings of a conference organized by the Institute of Mathematics and Its Applications on the mathematics of surfaces, held at Brunel University in September 1994. Clarendon Press, New York, NY, USA, 1996. ISBN 0-19-851198-1. LCCN QA571.M3853 1996.

Nguyen:2016:FEN

[NKP16] Thien Nguyen, Kȩstutis Karciauskas, and Jörg Peters. C^{1} finite elements on non-tensor-product 2 d and 3 d manifolds. Applied Mathematics and Computation, 272 (part 1)(??):148– 158, January 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S009630031500884X.

Ni:2007:TSQ

[NNP07] Tianyun Ni, Ahmad H. Nasri, and Jörg Peters. Ternary subdivision for quadrilateral meshes.

Computer-Aided Geometric Design, 24(6):361–370, 2007. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:1992:IHD

[NP92] A. Neff and J. Peters. C^{1} interpolation on higher-dimensional analogues of the 4-direction mesh. In Braess and Schumaker [BS92], pages 207–220. ISBN 3-7643-2746-4 (Basel), 0-8176-2746-4 (Boston). LCCN QA221 .N84 1992.

Nguyen:2016:NDG

[NP16a] Dang-Manh Nguyen and Jörg Peters. Nonuniform discontinuous Galerkin filters via shift and scale. SIAM Journal on Numerical Analysis, 54(3):1401– 1422, ???? 2016. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nguyen:2016:RSE

[NP16b] Thien Nguyen and Jörg Peters.

Refinable C^{1} spline elements for irregular quad layout. Computer-Aided Geometric Design, 43:123– 130, 2016. CODEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Nguyen:2017:ELD

[NP17] Dang-Manh Nguyen and Jörg Peters. Explicit least-degree boundary filters for discontinuous Galerkin. SIAM Journal on Scientific Computing, 39(4):A1741– A1765, ???? 2017. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Nairn:19xx:SQB

[NPLxx] D. Nairn, J. Peters, and D. Lutterkort. Sharp, quantitative bounds on the distance between a polynomial piece and its Bézier control polygon. Computer-Aided Geometric Design, ??(??), ?? 19xx. CODEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Ozkan:2018:AEA

[OPB^{+}18] Aysegul Ozkan, Rahul Prabhu, Troy Baker, James Pence, Jörg Peters, and Meera Sitharam. Algorithm 990: Efficient atlasing and search of configuration spaces of point-sets constrained by distance intervals. ACM Transactions on Mathematical Software, 44(4):48:1–48:30, August 2018. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic). URL https://dl.acm.
                                        org/citation.cfm?id=3204472.

Olshanskii:2006:UPP

[OPR06] Maxim A. Olshanskii, Jörg Peters, and Arnold Reusken. Uniform preconditioners for a parameter dependent saddle point problem with application to generalized Stokes interface equations. Numerische Mathematik, 105(1): 159–191, November 2006. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Peters:1989:LGH

[Pet89] Jörg Peters. Local generalized Hermite interpolation by quartic C^{2} space curves. ACM Transactions on Graphics, 8 (3):235–242, July 1989. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0730-
0301/77060.html.

Peters:1990:FSP

[Pet90a] J. Peters. Fitting smooth parametric surfaces to 3D data. PhD thesis, University of Wisconsin, Madison, WI, USA, 1990. PhD thesis; see also CMS Technical Report 91-2.

Peters:1990:LCB

[Pet90b] J. Peters. Local cubic and bicubic C^{1} surface interpolation with linearly varying boundary normal. Computer-Aided Geometric Design, 7:499–515, 1990. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:1990:LSS

[Pet90c] J. Peters. Local smooth surface interpolation: A classification. Computer-Aided Geometric Design, 7:191–195, 1990. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:1990:NSM

[Pet90d] J. Peters. The network simplex method on a multiprocessor. Networks, 20(7):845–859, 1990. CO-DEN NTWKAA. ISSN 0028-3045 (print), 1097-0037 (electronic).

Peters:1990:SMI

[Pet90e] J. Peters. Smooth mesh interpolation with cubic patches. Computer Aided Design, 22(2):109– 120, 1990. CODEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (electronic).

Peters:1991:PSE

[Pet91a] J. Peters. Parametrizing singularly to enclose vertices by a smooth parametric surface. In MacKay and Kidd [MK91], pages 1–7. ISBN 0-9695338-0-2. LCCN T385.P792 1991.

Peters:1991:SIM

[Pet91b] J. Peters. Smooth interpolation of a mesh of curves. Constructive Approximation, 7:221–247, 1991. ISSN 0176-4276 (print), 1432-0940 (electronic). Winner of SIAM Student Paper Competition 1989.

Peters:1992:IGS

[Pet92a] J. Peters. Improving G^{1} surface joins by using a composite patch. In Warren [War92], pages 345– 354. ISBN 0-8194-1031-4. LCCN T385 .C874 1992.

Peters:1992:JSP

[Pet92b] J. Peters. Joining smooth patches at a vertex to form a C^{k} surface. Computer-Aided Geometric Design, 9:387–411, 1992. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:1993:SFF

[Pet93] J. Peters. Smooth free-form surfaces over irregular meshes generalizing quadratic splines. Computer-Aided Geometric Design, 10:347–361, 1993. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:1994:CCM

[Pet94a] J. Peters. A characterization of connecting maps as roots of the identity. Curves and Surfaces in Geometric Design, pages 369–376, 1994.

Peters:1994:SAT

[Pet94b] J. Peters. Surfaces of arbitrary topology constructed from biquadratics and bicubics. In Sapidis [Sap94], pages 277–293. ISBN 0-89871-332-3. LCCN QA567 .D47 1994.

Peters:1994:EAE

[Pet94c] Jörg Peters. Evaluation and approximate evaluation of the multivariate Bernstein–Bézier form on a regularly partitioned simplex. ACM Transactions on Mathematical Software, 20(4): 460–480, December 1994. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/citations/journals/
toms/1994-20-4/p460-peters/.

Peters:1995:SBZb

[Pet95a] J. Peters. C^{2} surfaces built from zero sets of the 7-direction box spline. Mathematical Methods in Curve and Surface Design, 1995. ISBN ????

Peters:1995:BSS

[Pet95b] J. Peters. Biquartic C^{1} spline surfaces over irregular meshes. Computer Aided Design, 27(12), December 1995. CODEN CAIDA5.

ISSN 0010-4485 (print), 1879-2685 (electronic). To appear.

Peters:1995:SS

[Pet95c] Jörg Peters. C^{1}-surface splines. SIAM Journal on Numerical Analysis, 32(2):645–666, April 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peters:1995:BRB

[Pet95d] Jörg Peters. Book review: Geometric Concepts for Geometric Design (W. Boehm and H. Prautzsch). SIAM Review, 37(3): 473–474, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peters:1995:SPM

[Pet95e] Jörg Peters. Smoothing polyhedra made easy. ACM Transactions on Graphics, 14(2):162– 170, April 1995. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/toc/
                                        Abstracts/0730-0301/221670.
                                        html.

Peters:1995:SBZa

[Pet96a] J. Peters. C^{2} surfaces built from zero sets of the 7-direction box spline. In Mullineux [Mul96], pages 463–474. ISBN 0-19-851198-1. LCCN QA571.M3853 1996.

Peters:1996:CCS

[Pet96b] J. Peters. Curvature continuous spline surfaces over irregular meshes. Computer-Aided Geometric Design, 13(??), ???? 1996. CODEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic). To appear.

Peters:1996:IRC

[Pet96c] Jörg Peters. Interpolation regions for convex cubic curve segments. Advances in computational mathematics, 6(1):87–96, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/BF02127698.

Peters:1996:SSI

[Pet96d] Jörg Peters. Spline surfaces from irregular control meshes. Zeitschrift für Angewandte Mathematik und Mechanik, 76(1):69– 72, ?? 1996. CODEN ZAM-MAX. ISSN 0044-2267 (print), 1521-4001 (electronic). Proceedings of ICIAM ’95.

Peters:1997:CCM

[Pet97a] J. Peters. C code for modeling smooth free-form surfaces of arbitrary patch-layout with linearly-trimmed bicubic Bsplines (NURBS). ACM Transactions on Mathematical Software, ??(??), ???? 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). to appear.

Peters:1997:SPT

[Pet97b] J. Peters. Smoothing polyhedra using trimmed bicubic NURBS. In xxx. ????, ????, 1997. submitted.

Peters:1998:PSF

[Pet98a] J. Peters. Pcp2Nurb: smooth free-form surfacing with linearlytrimmed bicubic B-splines. ACM Transactions on Mathematical Software, x, 1998. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). to appear.

Peters:1998:SPU

[Pet98b] J. Peters. Smoothing polyhedra using trimmed bicubic patches. In Farin et al. [FBBD98], pages 211–216. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998.

                                        URL http://www.springer.at/
                                        springer.py?Page=30&Key=340&
                                        id=589&cat=3.

Peters:1998:APF

[Pet98c] Jörg Peters. Algorithm 783: Pcp2Nurb — smooth free-form surfacing with linearly trimmed bicubic B-splines. ACM Transactions on Mathematical Software, 24(3):261–267, September 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toms/1998-24-3/p261-
                                        peters/.

Peters:2001:SPR

[Pet01] Jörg Peters. Smooth patching of refined triangulations. ACM Transactions on Graphics, 20 (1):1–9, January 2001. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.acm.

org/pubs/citations/journals/
tog/2001-20-1/p1-peters/.

Peters:2002:FFS

[Pet02a] Jörg Peters. C^{2} free-form surfaces of degree (3, 5). Computer-Aided Geometric Design, 19(2):113– 126, 2002. CODEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:2002:GC

[Pet02b] Jörg Peters. Geometric continuity. In Handbook of computer aided geometric design, pages 193–227. North-Holland, Amsterdam, 2002.

Peters:2003:SFN

[Pet03] Jörg Peters. Smoothness, fairness and the need for better multisided patches. In Topics in algebraic geometry and geometric modeling, volume 334 of Contemp. Math., pages 55–64. Amer. Math. Soc., Providence, RI, 2003.

Peters:2004:MSL

[Pet04] Jörg Peters. Mid-structures linking curved and piecewise linear geometry. In Geometric modeling and computing: Seattle 2003, Mod. Methods Math., pages 453– 468. Nashboro Press, Brentwood, TN, 2004.

Peters:2012:CV

[Pet12] Jörg Peters. Changing variables. IEEE Computer Graphics and Applications, 32(3):88–93, May/June 2012. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Peters:2013:SUK

[Pet13] Jörg Peters. Splines and unsorted knot sequences. Computer-Aided Geometric Design, 30(7):733– 741, 2013. CODEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:2014:RSD

[Pet14] Jörg Peters. Refinability of splines derived from regular tessellations. Computer-Aided Geometric Design, 31(3-4):141–147, 2014. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:2015:GSF

[Pet15] Jörg Peters. General spline filters for discontinuous Galerkin solutions. Computers and Mathematics with Applications, 70(5):1046– 1050, 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Peters:2018:GSB

[Pet18] Jörg Peters. On G^{1} stitched bicubic Bézier patches with arbitrary topology. Computers and Graphics, 71(??):154–156, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https://www.

                                        sciencedirect.com/science/
                                        article/pii/S0097849317301966.

See response [ASC18].

Peters:2010:CSS

[PF10] Jörg Peters and Jianhua Fan. On the complexity of smooth spline surfaces from quad meshes.

Computer-Aided Geometric Design, 27(1):96–105, 2010. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:1994:GCC

[PH94] J. Peters and C. M. Hoffmann.

Geometric constraints of CAGD. In Dæhlen et al. [DLS94], page ???? ISBN 0-8265-1268-2. LCCN QA565 .M37 1995.

Peters:1998:PS

[PK98] J. Peters and L. Kobbelt. The platonic spheroids. Technical Report 97-052, Dept of Computer Sciences, Purdue University, 1998.

Peters:2023:ABC

[PLK23] Jörg Peters, Kyle Lo, and Kȩstutis Karciauskas. Algorithm 1032: Bi-cubic splines for polyhedral control nets. ACM Transactions on Mathematical Software, 49(1):7:1–7:??, March 2023. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL https://dl.acm.org/doi/
10.1145/3570158.

Peters:1997:CVS

[PN97] J. Peters and A. Nasri. Computing volumes of solids enclosed by recursive subdivision surfaces. Computer Graphics Forum, 16(3), September 1997. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peters:1997:SSS

[PR97] Jörg Peters and Ulrich Reif.

The simplest subdivision scheme for smoothing polyhedra. ACM Transactions on Graphics, 16(4): 420–431, October 1997. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

URL http://www.acm.org:80/
pubs/citations/journals/tog/
1997-16-4/p420-peters/.

Peters:1998:ECQ

[PR98a] J. Peters and U. Reif. The 42 equivalence classes of quadratic surfaces in affine n-space. Computer-Aided Geometric Design, 15, 1998. CODEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:1998:AAG

[PR98b] Jörg Peters and Ulrich Reif. Analysis of algorithms generalizing Bspline subdivision. SIAM Journal on Numerical Analysis, 35 (2):728–748, April 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30434.

Peters:2004:SCS

[PR04] Jörg Peters and Ulrich Reif. Shape characterization of subdivision surfaces—basic principles. Computer-Aided Geometric Design, 21(6):585–599, 2004. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).

Peters:2008:SS

[PR08] Jörg Peters and Ulrich Reif. Subdivision surfaces, volume 3 of Geometry and Computing. Springer-Verlag Inc., New York, NY, USA, 2008. ISBN 3-540-76405-4. xvi + 204 pp. With introductory contributions by Nira Dyn and Malcolm Sabin.

Peters:2005:FIS

[PRR05] Jörg Peters, Volker Reichelt, and Arnold Reusken. Fast iterative solvers for discrete Stokes equations. SIAM Journal on Scientific Computing, 27(2):646–666, March 2005. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/
article/60602.

Peters:1992:SVI

[PS92a] J. Peters and M. Sitharam. On stability of m-variate C^{1} interpolation. Approximation Theory and its Applications, 8(4):17–32, 1992. ISSN 1000-9221.

Peters:1992:SIC

[PS92b] Jörg Peters and Meera Sitharam.

Stability of interpolation from C^{1} cubics at the vertices of an underlying triangulation. SIAM Journal on Numerical Analysis, 29(2): 528–533, April 1992. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peters:2004:CDS

[PS04] Jörg Peters and Le-Jeng Shiue.

Combining 4- and 3-direction subdivision. ACM Transactions on Graphics, 23(4):980–1003, October 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peters:2015:PSS

[PS15] Jörg Peters and Martin Sarov. Polynomial spline surfaces with rational linear transitions. Computers and Graphics, 51(??): 43–51, October 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0097849315000606.

Peters:2006:EGR

[PSZF06] Jörg Peters, Meera Sitharam, Yong Zhou, and JianHua Fan. Elimination in generically rigid 3D geometric constraint systems. In Algebraic geometry and geometric modeling, Math. Vis., pages 205–216. Springer-Verlag Inc., New York, NY, USA, 2006.

Peters:2000:GMC

[PU00] Jörg Peters and Georg Umlauf. Gaussian and mean curvature of subdivision surfaces. In The mathematics of surfaces, IX (Cambridge, 2000), pages 59–69. Springer-Verlag Inc., New York, NY, USA, 2000.

Peters:2001:CCB

[PU01] Jörg Peters and Georg Umlauf. Computing curvature bounds for bounded curvature subdivision. Computer-Aided Geometric Design, 18(5):455–461, 2001. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic). Subdivision algorithms (Schloss Dagstuhl, 2000).

Peters:1997:BBI

[PW97a] J. Peters and M. Wittman. Blending basic implicit shapes using trivariate box splines. In Goodman and Martin [GM97], pages 409–426. ISBN 1-874728-12-7.

Peters:1997:BSB

[PW97b] J. Peters and M. Wittman. Boxspline based CSG blends. In W. Bronsvort C. Hoffmann, editor, Proceedings of Fourth Symposium on Solid Modeling. ACM Siggraph, May 1997.

Peters:2003:OPL

[PW03] Jörg Peters and Xiaobin Wu. On the optimality of piecewise linear max-norm enclosures based on slefes. In Curve and surface design (Saint-Malo, 2002), Mod. Methods Math., pages 335–344. Nashboro Press, Brentwood, TN, USA, 2003.

Peters:2006:LLI

[PW06] Jörg Peters and Xiaobin Wu.

On the local linear independence of generalized subdivision functions. SIAM Journal on Numerical Analysis, 44(6):2389– 2407, January 2006. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reif:2006:SAS

[RP06] Ulrich Reif and Jörg Peters.

Structural analysis of subdivision surfaces — a summary. In Topics in multivariate approximation and interpolation, volume 12 of Stud. Comput. Math., pages 149– 190. Elsevier, Amsterdam, The Netherlands, 2006.

Sapidis:1994:DFC

[Sap94] Nickolas S. Sapidis, editor. Designing fair curves and surfaces: shape quality in geometric modeling and computer-aided design. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-332-3. xii + 318 pp. LCCN QA567 .D47 1994.

Shiue:2005:RGS

[SJP05] Le-Jeng Shiue, Ian Jones, and Jörg Peters. A realtime GPU subdivision kernel. ACM Transactions on Graphics, 24(3):1010– 1015, July 2005. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sarov:2016:RPG

[SP16] Martin Sarov and Jörg Peters. Refinable polycube G-splines. Computers and Graphics, 58(??): 92–101, August 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0097849316300644.

Sitharam:2010:OPS

[SPZ10] Meera Sitharam, Jörg Peters, and Yong Zhou. Optimized parametrization of systems of incidences between rigid bodies. Journal of Symbolic Computation, 45(4):481–498, April 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sitharam:2010:RCC

[SZP10] Meera Sitharam, Yong Zhou, and Jörg Peters. Reconciling conflicting combinatorial preprocessors for geometric constraint systems. International Journal of Computational Geometry and Applications (IJCGA), 20(6):631–651, December 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Warren:1992:CSC

[War92] Joseph D. Warren, editor. Curves and surfaces in computer vision and graphics III: 16–18 November 1992, Boston, Massachusetts, volume 1830 of Proceedings of SPIE—the International Society for Optical Engineering. SPIE, International Society for Optical Engineering, Bellingham, WA, USA, 1992. ISBN 0-8194-1031-4. LCCN T385 .C874 1992.

Wu:2004:DFC

[WP04] Xiaobin Wu and Jörg Peters. Distance fields and collision detection: Interference detection for subdivision surfaces. Computer Graphics Forum, 23(3):577–584, September 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yeo:2014:EPA

[YBP14] Young In Yeo, Sagar Bhandare, and Jörg Peters. Efficient pixelaccurate rendering of animated curved surfaces. In Mathematical methods for curves and surfaces, volume 8177 of Lecture Notes in Computer Science, pages 491–509.

Springer-Verlag Inc., New York, NY, USA, 2014.

Yang:2018:CGN

[YP18] Huogen Yang and Jörg Peters. Constraints for geodesic network interpolation at a vertex. Computer-Aided Geometric Design, 67:71–78, 2018. CO-DEN CAGDEX. ISSN 0167-8396 (print), 1879-2332 (electronic).